what is the probability that you will get heads four times is a row when flipping a fair coin?
a.0.5
b.0.25
c.0.125
d.0.0625

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

what is the probability that you will get heads four times is a row when flipping a fair coin?
--------------------
Ans: (1/2)^4 = 1/16 = 0.0625
----------------
Cheers,
Stan H.
----------------
a.0.5
b.0.25
c.0.125
d.0.0625

Answer by rfer(16322)

1/2*1/2*1/2*1/2=1/16=(D 0.0625)